McGilla Golf is evaluating a new golf club. The clubs will sell for $780 per set and have a variable cost of $380 per set. The company has spent $148,000 for a marketing study that determined the company will sell 52,000 sets per year for seven years. The marketing study also determined that the company will lose sales of 9,300 sets of its high-priced clubs. The high-priced clubs sell at $1,080 and have variable costs of $680. The company will also increase sales of its cheap clubs by 10,800 sets. The cheap clubs sell for $420 and have variable costs of $220 per set. The fixed costs each year will be $9,080,000. The company has also spent $1,090,000 on research and development for the new clubs. The plant and equipment required will cost $28,560,000 and will be depreciated on a straight-line basis. The new clubs will also require an increase in net working capital of $1,280,000 that will be returned at the end of the project. The tax rate is 36 percent, and the cost of capital is 10 percent. What is the sensitivity of the NPV to the price and quantity of the new clubs? (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

NPV

ΔNPV/ΔP $

ΔNPV/ΔQ $
Answer #1

1.
=52000*1*(1-36%)/10%*(1-1/1.1^7)=162020.978252821

2.
=1*(780-380)*(1-36%)/10%*(1-1/1.1^7)=1246.31521732939